APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Wednesday March 11, 2009 (MJD 54902) ---=== OBSERVING TEAM ===--- Elena Malanushenko Kaike Pan Viktor Malunushenko (swing) Marvels team: Scot Fleming (UF) Lauren Palladino (Vanderbilt Univ) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- The first 2.5 hours was lost to bad weather. After open we acquired signals on plates: 3089B, 3277B, 3280A, 3280B, 3359A, 3359B. Problem: 1) There is a problem in guider fiber #4 for plate 3089 B. likely, the problem is in plate design file. (see log for detail). 2) the 4 clock FF lamp was burnt out. The bad one was put on Wendell's office desk. ---=== OBSERVING LOG ===--- Afternoon: ---------- Filesystem Size Used Avail Use% Mounted on /dev/sda3 142G 116G 20G 86% / /dev/sda1 99M 30M 64M 33% /boot /dev/sdb1 917G 801G 70G 92% /data sdssfiles1p:/data1/data 1.5T 671G 722G 49% /files All systems were checked and working properly. The spectro focus was checked but no adjustment made. The collimate is set for about +4.5 C. We were plan to start ventilation at 00:00Z, but the humidity changed so quickly(dewpoint spread could bounce around from 5 to 2 within 10 mintutes) we decided to not. Night: ------ 01:00Z mounted cart 7, for a marvel test. 03:15Z humidity dropped. We started ventilation and mounted cartridge 1. 04:35Z enclousure off. 04:50Z MARVELS calibration when filling SDSS dewars. 05:05Z cart 1, 3089B. SOGTF. 05:15Z, start a 50-min exposure. Guider fiber #4 does not work properly. When the guider was started, there is no measurement for this fiber although we do see star image there. Then Elena issued an command "enableFiber 4", SoP thraw an error: Cannot read Guider Data (4.84, efficency): no such element in the array, and guider stoped to work. She had to do stop guider, disable the fiber, and restart the guider. 06:10Z Cart 5, 3277A. One of FF lamps, the 4-clock one, burnt out when slewing to the field. So no guider flat was taken, we had to change it. The bad one is one wendell's office desk. About 20 mointues was lost, since we are runing before shedule, Marvels team decided to skip this field, and do the next. If count this swap change, goto field time, 35 mintues was lost due to the lamp problem. 06:40Z Marvels calibration. 06:50Z cart 4, 3280A. FK5 used. one 50 mintues exposure was taken on this field. 07:50Z calibration. 08:00Z cart 4, 3280B. SOGTF. 50 min exposure and a bracket of calibration. 09:05Z cart 5, 3277B. SOGTF. the 1st guider frame is 1205. 50 min exposure and a bracket calibration. 10:10Z cart 8, plate 3359A. SOGTF. one 50-min exposure and a bracket calibration. 11:20Z cart 8, plate 3359B. SOGTF. one 50-min exposure and a bracket calibration. 12:40Z endNight rejoiced. Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 59647, plate 3231 ------- 23:36 59647 10.0 flat 23:39 59648 2.0 arc 23:46 59649 3.0 arc 23:49 59650 3.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 05:42Z 1 3089 224.33 0.0007 60.42 0.0017 66.57 0.0288 0.999890 07:11Z 4 3280 -42.43 0.0007 64.18 0.0013 144.47 0.0131 0.999850 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 05:42Z 1 3089 0.99989 1085 2100 50 224 60.4 -0.3 7 299 -- 2.0 07:12Z 4 3280 0.99985 1480 2468 100 -43 64.2 -0.7 2 226 -- 1.6 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:10Z 38 28 7 242 (WSW) 98 - clear 23:40Z 37 28 5 263 (W) 137 - 00:10Z 36 30 5 263 (W) 149 - 00:46Z 34 31 2 193 (SSW) 227 - 01:17Z 31 30 1 114 (ESE) 238 - 01:47Z 31 29 1 250 (WSW) 318 - 02:19Z 31 28 3 306 (NW) 278 - 02:51Z 31 26 4 274 (W) 311 - 03:23Z 31 23 5 291 (WNW) 395 - clear 03:55Z 30 25 2 254 (WSW) 370 - 04:27Z 30 23 4 339 (NNW) 436 - 05:00Z 30 20 0 65 (ENE) 495 - 05:32Z 31 20 3 328 (NNW) 499 - 06:04Z 31 21 5 269 (W) 435 - 06:36Z 31 21 2 266 (W) 423 - 07:08Z 30 24 2 239 (WSW) 454 - 07:41Z 30 24 3 252 (WSW) 409 - 08:13Z 31 20 5 261 (W) 467 - 08:45Z 32 20 6 263 (W) 403 - 09:17Z 31 21 6 264 (W) 375 - 09:49Z 31 21 7 303 (WNW) 322 - 10:20Z 31 21 7 317 (NW) 329 - 10:52Z 31 22 4 260 (W) 375 - 11:22Z 29 22 5 263 (W) 421 - mostly clear 11:54Z 30 23 8 248 (WSW) 237 - 12:26Z 30 22 7 225 (SW) 356 - most clear ---=== TELESCOPE STATUS ===--- 3:15Z - Began ventilation. Doors and louvers opened, fans on. 4:35Z - Enclosure off. 12:35Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at: 12:45Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 8 Counterweights at: 255 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 151 lbs, 23 psi Imager lbs, psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_39_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.10.0 2009-01-12 sdssProcedures: SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0310 -0.0280 Axial B 0.0270 0.0470 Axial C -0.0370 -0.0360 Trans D -0.7620 -0.7600 Lateral E 10.6045 10.5791 Lateral F 12.1793 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. 10350. 10000. Actual: 10471. -76. 5755. 8616. 10374. 10023. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 495.77 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: 2.6 Alt.: 30.068733 MIGS TONIGHT NOMINAL Axial A 18.5870 18.5840 Axial B 15.6300 15.6220 Axial C 13.8160 13.8150 Trans D 15.1040 15.1080 GALILS Commanded: 1626666. 1528733. 1595491. -2950. -4550. Actual: 1626627. 1528808. 1595583. -2945. -4561. SETMIR VALUES SecDesOrient: 1257.00 -12.00 -28.00 -39.00 94.77 SecOrient: 1256.97 -11.98 -28.04 -39.21 94.69 ---=== PROBLEMS IN DETAIL ===--- Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00059646.fit -5.5 -3.9 -2.3 2.7 7.1 9.2 11.3 sdR-b2-00059646.fit -8.0 -5.5 -4.3 1.9 7.2 9.3 11.0 sdR-r1-00059646.fit -6.5 -4.9 -3.5 2.5 8.2 9.5 12.5 sdR-r2-00059646.fit -7.8 -5.7 -4.0 1.6 7.9 10.0 13.2 ------------------- ------ ------ ------ ------ ------ ------ ------ sp1> Coll_motor_A -5777 01P-0000005777 Coll_motor_B -5950 02P-0000005950 Coll_motor_C -5881 01P-0000005881 Temp_Median 6.9 sp2> Coll_motor_A -4193 01P-0000004193 Coll_motor_B -4743 02P-0000004743 Coll_motor_C -4893 01P-0000004893 Temp_Median 5.7 --------------------------------- IDL 5.5a> collimate, 59649 % Compiled module: COLLIMATE. Reading /data/spectro/54902/sdR-b1-00059649.fit Reading /data/spectro/54902/sdR-b1-00059650.fit COLLIMATE: Min offset = -0.69 pix COLLIMATE: Max offset = 1.03 pix COLLIMATE: RMS across CCD = 0.44 pix COLLIMATE: Mean offset = 0.21 pix COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -7.6 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-54902-b1-00059649.log PostScript file = Collimate-54902-b1-00059649.ps Reading /data/spectro/54902/sdR-b2-00059649.fit Reading /data/spectro/54902/sdR-b2-00059650.fit COLLIMATE: Min offset = -0.22 pix COLLIMATE: Max offset = 0.13 pix COLLIMATE: RMS across CCD = 0.08 pix COLLIMATE: Mean offset = 0.01 pix COLLIMATE: Camera b2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of -0.5 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-54902-b2-00059649.log PostScript file = Collimate-54902-b2-00059649.ps Reading /data/spectro/54902/sdR-r1-00059649.fit Reading /data/spectro/54902/sdR-r1-00059650.fit COLLIMATE: Min offset = -0.62 pix COLLIMATE: Max offset = 0.22 pix COLLIMATE: RMS across CCD = 0.25 pix COLLIMATE: Mean offset = -0.24 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 2230 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 2230" Log file = Collimate-54902-r1-00059649.log PostScript file = Collimate-54902-r1-00059649.ps Reading /data/spectro/54902/sdR-r2-00059649.fit Reading /data/spectro/54902/sdR-r2-00059650.fit COLLIMATE: Min offset = -0.79 pix COLLIMATE: Max offset = 0.15 pix COLLIMATE: RMS across CCD = 0.26 pix COLLIMATE: Mean offset = -0.12 pix COLLIMATE: Camera r2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 1119 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 1119" Log file = Collimate-54902-r2-00059649.log PostScript file = Collimate-54902-r2-00059649.ps